what is the answer to 11^4/11^2? show steps.
Answer by richwmiller(17219)

Where are the negative or fractional exponents?
two ways to solve
subtracting exponents 4-2=2
11^2=121
and
11*11*11*11/!!*11
=11*11=11^2=121